OpenStudy (blake3299):

WHAT IS THE SUM OF 8/17 +5/17 A. 13/289 B. 40/289 C. 13/34 D. 13/17

OpenStudy (blake3299):

I SAY C

TheSmartOne (thesmartone):

\(\sf\Large\frac{a}{c}+\frac{b}{c}=\frac{a+b}{c}\)

TheSmartOne (thesmartone):

Your answer is wrong.

OpenStudy (anonymous):

it is D

TheSmartOne (thesmartone):

@joannacaro123 Do not give direct answers as it is against the Code of Conduct. http://openstudy.com/code-of-conduct

OpenStudy (bee_see):

You just have to add the numerators since the denominator is the same for both fractions.

OpenStudy (anonymous):

so 8+7 =15 so you have 15/17. Then you need to multiply the denominator by a number to get the denominator that is one of the choices. Then you do what you did to the numerator Does this help?

OpenStudy (anonymous):

oh thanks i didnt know sorry @TheSmartOne